1. Something is wrong with the student’s work below.
Explain what the student did wrong.
Next, show how to simplify the fraction correctly.
The problem : x^2+6x+9 / x+3
Sara simplifies the expression as follows:
x^2+6(x is crossed out and replaced with ^1)+(9 is crossed out and replaced with 3) / (x is replaced with 1)+(3 is replaced with 1). It should look like this, x^2+6^1+3 / 1+1 = x^2+9, (x=-3)
Explain what the student did wrong.
Next, show how to simplify the fraction correctly.

Answer by richwmiller(17219)

Did this a few days ago.
(x^2+6x+9) / (x+3)=x+3
(x+3)^2=x^2+6x+9